I please need help on ways to remember how to multiply, subtract, add, divide negative and positive numbers. Please, ive tried all the ways ive been taught.
Thank you a lot

Answer by apachemexican1(64)

ok when adding and subtracting negative numbers
is very simple.
when adding and subtracting
the bigger number is gonna carry the sign
if the bigger number is negative
its gonna be negative
if the bigger number is positive
its gonna be positive
ex: -12+4=
since -12 is bigger the answer is gonna be negative
-12+4= -8
ex: 12-4=
since 12 is bigger number the answer will be positive.
12-4=8
*Now the way I remember about multiplying and dividing
negative and positive numbers is really cool.
Its called the love boat!
*now if two people are in a reltionship are
always positive then their re;ationship will be positive
*now if two people are in a relationship and they are
always negative, they agree on everything then their
relationship will be positive.
*now if two people are in a relatioship, one is always
negative and the other is positive, they will never
agree so the relationship will be negative
I hope you understand this.
here are a few examples.
2*2= 4 (two positives equal a positive number)
-2*2=-4 (one negative one positive equals a negative number)
-2*-2=4 (two negative numbers equals a positive number)
good luck!